OpenStudy (anonymous):

Graph y = cosx for - ≤ x ≤ -. What is the largest value in the range? -1 0 1

myininaya (myininaya):

do you know the unit circle?

OpenStudy (anonymous):

yes

myininaya (myininaya):

so on the unit circle, what values does the x-value stay between?

myininaya (myininaya):

or you don't know that do you believe in any right triangle the hyp is always the biggest?

OpenStudy (anonymous):

idk im on odyssyware the unit circle not on here

OpenStudy (anonymous):

wouldnt the answer be 1 ?

myininaya (myininaya):

yes the range of both y=sin(x) and y=cos(x) is [-1,1] These values on the unit circle for x and y stay between -1 and 1 (inclusion) so the max value for either is 1

OpenStudy (anonymous):

thank youuu so much

OpenStudy (anonymous):

Which of the following ordered pairs lies on the graph of y = tanx? (-, -1) (-, 1) (, ) (5π, 0)

OpenStudy (anonymous):

Which of the following ordered pairs lies on the graph of y = tanx? (-, -1) (-, 1) (, ) (5π, 0)

myininaya (myininaya):

I only see one pair of numbers so I would try to see if that one pair works.

myininaya (myininaya):

\[0=\tan(5 \pi)=\frac{\sin(5\pi)}{\cos(5\pi)}\]

myininaya (myininaya):

you want to see if that equality holds

myininaya (myininaya):

if it holds then (5pi,0) is a point on y=tan(x)

myininaya (myininaya):

and I guess your question before you were actually given the range you just didn't type any of the numbers ?